Write the following in terms of cos x only, if the answer needs a ± don't list both functions, just answer the multiple choice portion appropriately


1
Expert's answer
2021-09-14T06:05:09-0400

"1. \ sin^2x+cos^2x=1\\
\Rightarrow sin^2x=1-cos^2x\\
\Rightarrow sinx=\pm\sqrt{1-cos^2x}\\
\\
2. \ tanx=\frac{sinx}{cosx} \\
\Rightarrow tanx=\pm\frac{\sqrt{1-cos^2x}}{cosx}\\
3. \ secx =\frac{1}{cosx}\\
4. \ cosecx=\frac{1}{sinx}=\pm\frac{1}{\sqrt{1-cos^2x}}\\
5.\ cotx=\frac{1}{tanx}=\frac{cosx}{sinx}=\pm\frac{cosx}{\sqrt{1-cos^2x}}"